OpenStudy (anonymous):

Inverse of Log base 4 of x

OpenStudy (anonymous):

y=4^x\[f(x)=4^x\] ?

OpenStudy (anonymous):

Yes Perto! xD

OpenStudy (anonymous):

*gasp i was thinking that! well there's one i got right on the test..

OpenStudy (anonymous):

Now, try this one! Log base 2 of x + 3

OpenStudy (anonymous):

the inverse of that?

OpenStudy (anonymous):

yesh!

OpenStudy (anonymous):

f(x)= 2^(x+3)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

f(x)= 2^x ?

OpenStudy (anonymous):

wait a sec!

OpenStudy (anonymous):

wait lemme see!

OpenStudy (anonymous):

where's hanh?

OpenStudy (anonymous):

The answer is y = 2^x - 3

OpenStudy (anonymous):

why minus?

OpenStudy (anonymous):

can you show me how you did that

OpenStudy (anonymous):

cuz its the inverse

OpenStudy (anonymous):

darn..that means i probably got that wrong in the test. :\

OpenStudy (anonymous):

Okie! So it's: y = Log base e of x + 3, NOT (x+3) So it's 2^y = x + 3 There fore, 2^x - 3 = y

OpenStudy (anonymous):

id remember if i got that right

OpenStudy (anonymous):

base 2*

OpenStudy (anonymous):

next problem

OpenStudy (anonymous):

Log base x of 3x

OpenStudy (anonymous):

thats wierd

OpenStudy (anonymous):

x ^ y = 3x Inverse = x^x = 3y, so it's y = (x^x) / 3

OpenStudy (anonymous):

thats wierd

OpenStudy (anonymous):

okay, I'll change base x to base 2

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

so that you won't be confuse with which x to switch

OpenStudy (anonymous):

2^y or x^y in the original = 3x Inverse = 2^x original x^x = 3y so divide both side by 3 to get y y = (2^x)/3 or (x^x)/3 in the original